x - square root of x = 0, solve for all values of x that satisfies the equation show work
Answer by stanbon(75887)

x - square root of x = 0,
Add sqrt(x) to both sides
sqrt(x) = x
Square both sides to get:
x=x^2
x^2-x=0
x(x-1)=0
x=0 or x=1
Both are good answers.
-------
Cheers,
Stan H.